# Management and Visualization Tools for Emergency Medical Services

### v1

@@ Abstract
anton@isye.gatech.edu
@@
UFRJ
Rio de Janeiro, Brazil
@@ 1 Introduction
The management of an emergency medical service (EMS) is a complex process that includes the real-time allocation of ambulances and crews to emergency calls received at a call center. Tools for the visualization of emergency data, forecasts, ambulance movements, and performance metrics, are of great benefit in the planning and real-time control of emergency services, the training of personnel, and the development of management tools. In this paper, we describe a visualization tool that we developed for the efficient management of an EMS. The tool is available at http://samu.vincentgyg.com/. The tool also includes statistical methods for forecasting medical emergencies described in [22] and [23], available at https://github.com/vguigues/LASPATED, as well as ambulance dispatch algorithms described in [21], and available at https://github.com/vguigues/Heuristics_Dynamic_Ambulance_Management.
@@
Research on ambulance management has addressed various planning and operational problems. Many papers have addressed the location of ambulance stations or the assignment of ambulances to stations, including [56, 5, 10, 49, 13, 12, 28, 48, 46, 17, 29, 15, 51], and [52]. Various stochastic models, including queuing models and simulations, have been proposed by [59, 54, 55, 16, 35, 36, 27, 32, 18, 19, 20, 9] and [47] for evaluating location decisions for stations and ambulances. The most popular policy in the literature for the ambulance selection problem is the simple closest-available-ambulance rule, used by [25, 26, 40, 42, 41], and [1]. As the name indicates, when an emergency call arrives, the closest-available-ambulance rule dispatches the available ambulance that is closest (in terms of forecasted response time) to the emergency. A few papers have proposed alternatives to the closest-available-ambulance rule, including [2, 37, 50, 3, 43, 4, 38, 30] and [31]. Fewer papers have addressed the ambulance reassignment decision (where to send an ambulance when it becomes available. One popular approach is the following: First, the ambulance stations are located. Second, each ambulance is assigned to a station, called its home station. Then, during operations, when an ambulance becomes available and is not dispatched to an emergency waiting in queue, the ambulance is sent to its home station [18, 26, 47, 3, 33, 39, 43, 4]. In [24], a stochastic optimization model is proposed for both the ambulance selection problem and the ambulance reassignment problem.
@@ 4.1 Simulation of ambulance operations
[…]
@@
[…]
@@
[…]
@@ 4.2 Performance metrics
[…]
@@ 4.3 Ambulance selection and reassignment heuristics
In this section, we provide a brief description of the ambulance selection and ambulance reassignment heuristics available through the webpage http://samu.vincentgyg.com/ (the paper [21] contains more details and performance comparisons for these heuristics). These heuristics are the Closest Available (CA) heuristic, Best Myopic (BM) heuristic, a NonMyopic (NM) heuristic, GHP1 (Greedy Heuristic with Priorities 1), and GHP2 (Greedy Heuristic with Priorities 2).
@@
NM heuristic.
The NM heuristic assumes that we know, potentially, all the calls that will arrive in a given future time window, but we do not send ambulances to calls before these calls actually arrive. Under uncertainty in the time and location of calls, it can be used in a rolling horizon approach to solve second stage scenarios as in Section 3 of [21]. For each call $i$ for which no ambulance has been sent, NM computes the set $\mathcal{S}(i)$ of best ambulances which are the ambulances that achieve the minimal allocation cost, given previous ambulance allocations, where the allocation cost is given by (3). If any ambulances in $\mathcal{S}(i)$ are available, then NM dispatches a least advanced ambulance (as basic as possible) from $\mathcal{S}(i)$. Otherwise, we compute for each ambulance $j\in\mathcal{S}(i)$, the set $\mathcal{T}(j)$ of calls that arrive not later than $t_{f}(j)$ and such that $j$ is one of the best ambulances for that call, in terms of allocation cost. An ambulance $j\in\mathcal{S}(i)$ is termed good for call $i$ if, for every call $k\in T(j)$, the cost of allocating $j$ to $i$ is less than or equal to the cost of allocating $j$ to $k$. If the ambulance $j$ is good for call $i$, we dispatch $j$ to $i$ after it finishes its current service. Otherwise, we dispatch $j$ to the call $k\in T(j)$ that achieves the minimal allocation cost. When an ambulance after service is not immediately sent to an emergency, it is sent to its home base (if the home base rule is used) or to the closest station.
@@
In GHP1, every time a call arrives, it is put in the queue of calls, possibly for immediate dispatch, see below. Given the current time $t$ (this is a time at which we need to make a decision: when a call arrives or when an ambulance finishes service), we sort the queue of calls by decreasing order of the penalized waiting time. Specifically, for call $i$ of type $\mbox{Type}_{c}(i)$ that arrived at time $t_{c}(i)$, the penalized waiting time is penalization$(t-t_{c}(i),\mbox{Type}_{c}(i))$. Considering the calls in decreasing order of the penalized waiting time, we do for every call the following: we compute the ambulances that achieve the smallest allocation cost for that call. If any of them are available, we send to that call, among these ambulances, an ambulance as basic as possible. Otherwise the call remains in the queue of calls, and we proceed to the next call in queue. When an ambulance is sent to a station, it goes either to the closest station or to its home base.
@@ 4.5 Visualization of ambulance movements
The user can also choose the time acceleration factor: ×$\times 2$, ×$\times 10$, ×$\times 50$, ×$\times 100$, ×$\times 200$, or ×$\times 400$. For example, the acceleration factor ×$\times 100$ runs a simulation for a time window of $120$ minutes in $120/100$ minutes or $72$ seconds of user time. Ambulance trajectories are updated every $t_{\mathrm{step}}$ seconds of simulated time using the computations described in Section 4.6. Thus, with acceleration factor ×$\times 100$, ambulance trajectories are updated every $t_{\mathrm{step}}/100$ seconds of user time. A legend is given on the map, which shows $12$ types of information:
@@ 4.6 Computation of ambulance trajectories
Recall that latitudes are between $-90^{\circ}$ and $90^{\circ}$ while longitudes are between $-180^{\circ}$ and $180^{\circ}$. The Cartesian coordinates of $P_{1}$ and $P_{2}$ are given by
@@
for $i=1,2$, where $R=6371$km is the average radius of the earth. The time for the ambulance to go from $P_{1}$ to $P_{2}$ at constant speed $v$ is $\Delta=2R\arcsin(\|P_{2}-P_{1}\|/2R)/v$. Therefore, if $t\geq t_{\mathrm{prev}}+\Delta$, then the ambulance arrives at destination $P_{2}$ no later than $t$. If $t<t_{\mathrm{prev}}+\Delta$, then to compute the latitude and longitude of $P$, we refer to Figure 13, where we have represented $P_{1}$, $P_{2}$, $P$, the intersection $Q$ between $[P_{1},P_{2}]$ and $[O,P]$ (O is the center of the earth), angle $\alpha=\angle P_{1}OP_{2}=2\arcsin(\|P_{2}-P_{1}\|/2R)$, and $\alpha_{0}=\angle P_{1}OP=v(t-t_{\mathrm{prev}})/R$.
Figure 13: Bottom plot: geodesic from $P_{1}$ to $P_{2}$ with $\alpha=\angle P_{1}OP_{2}$ and point $P$ on this geodesic with $\alpha_{0}=\angle P_{1}OP$. On the 2-D top plot, we also represented the intersection $Q$ of $[O,P]$ and $[P_{1},P_{2}]$
@@
$OQ\ \ =\ \ \frac{P_{1}P_{2}\cos(\alpha/2)}{\sin(\alpha-\alpha_{0})+\sin(\alpha_{0})}\ \ =\ \ \frac{2R\sin(\alpha/2)\cos(\alpha/2)}{\sin(\alpha-\alpha_{0})+\sin(\alpha_{0})}\ \ =\ \ \frac{R\sin(\alpha)}{\sin(\alpha-\alpha_{0})+\sin(\alpha_{0})}.$ (9)
@@
3:$t_{\mathrm{next}}\leftarrow\mbox{Times}(j)(2)$
4:$t_{\mathrm{prev}}=\mbox{Times}(j)(1)$
5:$t_{\mathrm{current}}\leftarrow t_{\mathrm{prev}}-(t_{\mathrm{prev}}\mod t_{\mathrm{step}})$
6:if $t_{\mathrm{current}}=t_{\mathrm{prev}}$ then
@@ 4.8 Running the ambulance dispatch heuristics with code available on GitHub
The subdirectory SAMU3 of the GitHub repository contains the implementations of the heuristics described in section 4.3, considering the graph of the streets of Rio de Janeiro. The implementations can be found in the source file src/solver.cpp. A class was built for every heuristic. These classes are QueueSolver for CA heuristics, ForwardSolver for BM heuristic, NonMyopicSolver for NM heuristic, PrioritySolver for GHP1 heuristic, and MinMaxSolver for GHP2 heuristic. The code is written in C++ and has the following dependencies:

### v2

@@ Abstract
anton@isye.gatech.ed
@@
School of Applied Mathematics, FGV
Praia de Botafogo, Rio de Janeiro, Brazil
@@ 1 Introduction
The management of an emergency medical service (EMS) is a complex process that includes the real-time allocation of ambulances and crews to emergency calls received at a call center. Tools for the visualization of emergency data, forecasts, ambulance movements, and performance metrics, are of great benefit in the planning and real-time control of emergency services, the training of personnel, and the development of management tools. In this paper, we describe a visualization tool that we developed for the efficient management of an EMS. The tool is available at http://samu.vincentguigues.com. The tool also includes statistical methods for forecasting medical emergencies described in [22] and [23], available at https://github.com/vguigues/LASPATED, as well as ambulance dispatch algorithms described in [21], and available at https://github.com/vguigues/Heuristics_Dynamic_Ambulance_Management.
@@
Research on ambulance management has addressed various planning and operational problems. Many papers have addressed the location of ambulance stations or the assignment of ambulances to stations, including [56, 5, 10, 49, 13, 12, 28, 48, 46, 17, 29, 15, 51], and [52]. Various stochastic models, including queuing models and simulations, have been proposed by [59, 54, 55, 16, 35, 36, 27, 32, 18, 19, 20, 9] and [47] for evaluating location decisions for stations and ambulances. The most popular policy in the literature for the ambulance selection problem is the simple closest-available-ambulance rule, used by [25, 26, 40, 42, 41], and [1]. As the name indicates, when an emergency call arrives, the closest-available-ambulance rule dispatches the available ambulance that is closest (in terms of forecasted response time) to the emergency. A few papers have proposed alternatives to the closest-available-ambulance rule, including [2, 37, 50, 3, 43, 4, 38, 30] and [31]. Fewer papers have addressed the ambulance reassignment decision (where to send an ambulance when it becomes available). One popular approach is the following: First, the ambulance stations are located. Second, each ambulance is assigned to a station, called its home station. Then, during operations, when an ambulance becomes available and is not dispatched to an emergency waiting in queue, the ambulance is sent to its home station [18, 26, 47, 3, 33, 39, 43, 4]. In [24], a stochastic optimization model is proposed for both the ambulance selection problem and the ambulance reassignment problem.
@@ 4.1 Simulation of ambulance operations
  • the penalty $\mbox{WaitingOnScenePenalized}(i)$ for the amount of time between the instant of the emergency call and the instant that an ambulance arrives on the scene of the emergency: the penalized waiting time is
    $\mbox{WaitingOnScenePenalized}(i)\\ ={\tt penalization}(\mbox{WaitingOnScene}(i),\mbox{Type}_{c}(i))$
@@
  • the amount of time $\mbox{WaitingToHospital}(i)$ between the instant that the ambulance arrives on the scene of the emergency and the instant that the ambulance arrives at the hospital, if applicable;
  • the penalty $\mbox{WaitingToHospitalPenalized}(i)$ is given by
@@
After ambulance indexAmb leaves the scene of emergency $i$, it travels to a hospital at location $\ell_{h}(i)$. Thus, $\mbox{TripType}(\mbox{indexAmb})(j+3)=4$, $\mbox{AmbulancesTrips}(\mbox{indexAmb})(j+4)=\ell_{h}(i)$, and $\mbox{AmbulancesTimes}(\mbox{indexAmb})(j+4)=\mbox{AmbulancesTimes}(\mbox{indexAmb})(j+3)+\mbox{travelTime}(\ell_{c}(i),\ell_{h}(i),\mbox{AmbulancesTimes}(\mbox{indexAmb})(j+3))$. Also, $\mbox{WaitingToHospital}(i)=\mbox{TimeOnScene}(i)+\mbox{travelTime}(\ell_{c}(i),\ell_{h}(i),\mbox{AmbulancesTimes}(\mbox{indexAmb})(j+3))$.
@@ 4.2 Performance metrics
$\mbox{cost\_allocation\_ambulance}(a,c,t)\ \ =\ \ \mbox{{\tt{penalization}}}(t,c)+M_{ac}$ (3)
@@ 4.3 Ambulance selection and reassignment heuristics
In this section, we provide a brief description of the ambulance selection and ambulance reassignment heuristics available through the webpage http://samu.vincentguigues.com/ (the paper [21] contains more details and performance comparisons for these heuristics). These heuristics are the Closest Available (CA) heuristic, Best Myopic (BM) heuristic, GHP1 (Greedy Heuristic with Priorities 1), and GHP2 (Greedy Heuristic with Priorities 2).
@@
[…]
@@ 4.5 Visualization of ambulance movements
The user can also choose the time acceleration factor: $\times 2$, $\times 10$, $\times 50$, $\times 100$, $\times 200$, or $\times 400$. For example, the acceleration factor $\times 100$ runs a simulation for a time window of $120$ minutes in $120/100$ minutes or $72$ seconds of user time. Ambulance trajectories are updated every $t_{\mathrm{step}}$ seconds of simulated time using the computations described in Section 4.6. Thus, with acceleration factor $\times 100$, ambulance trajectories are updated every $t_{\mathrm{step}}/100$ seconds of user time. A legend is given on the map, which shows $12$ types of information:
@@ 4.6 Computation of ambulance trajectories
Recall that latitudes are between $90^{\circ}$ and $90^{\circ}$ while longitudes are between $180^{\circ}$ and $180^{\circ}$. The Cartesian coordinates of $P_{1}$ and $P_{2}$ are given by
@@
for $i=1,2$, where R=6371R=6371km is the average radius of the earth. The time for the ambulance to go from $P_{1}$ to $P_{2}$ at constant speed $v$ is $\Delta=2R\arcsin(\|P_{2}-P_{1}\|/2R)/v$. Therefore, if $t\geq t_{\mathrm{prev}}+\Delta$, then the ambulance arrives at destination $P_{2}$ no later than $t$. If $t<t_{\mathrm{prev}}+\Delta$, then to compute the latitude and longitude of $P$, we refer to Figure 13, where we have represented $P_{1}$, $_{2}, PP$, the intersection $Q$ between $P_{1},P_{2}]$ and $[O,P]$ (O is the center of the earth), angle $\alpha=\angle P_{1}OP_{2}=2\arcsin(\|P_{2}-P_{1}\|/2R)$, and $\alpha_{0}=\angle P_{1}OP=v(t-t_{\mathrm{prev}})/R$.
Figure 13: Bottom plot: geodesic from $P_{1}$ to $P_{2}$ with $\alpha=\angle P_{1}OP_{2}$ and point $P$ on this geodesic with $\alpha_{0}=\angle P_{1}OP$. On the 2-D top plot, we also represented the intersection $Q$ of $[O,P]$ and $P_{1},P_{2}]$
@@
[…]
@@
[…]
6:if $t_{\mathrm{current}}=t_{\mathrm{prev}}$ then
7:  discretized_rides(j) $\leftarrow$ {Trips(j)(1)}
8:  discretized_ride_types(j) $\leftarrow$ {Types(j)(1)}
@@ 4.8 Running the ambulance dispatch heuristics with code available on GitHub
The subdirectory SAMU3 of the GitHub repository contains the implementations of the heuristics described in section 4.3, considering the graph of the streets of Rio de Janeiro. The implementations can be found in the source file src/solver.cpp. A class was built for every heuristic. These classes are QueueSolver for CA heuristics, ForwardSolver for BM heuristic, PrioritySolver for GHP1 heuristic, and MinMaxSolver for GHP2 heuristic. The code is written in C++ and has the following dependencies: